A 1610kg car accelerates from zero to 100km/h in 5.50 s. Calculate it’s average power output of the engine
1
Expert's answer
2019-05-13T11:30:17-0400

Initially the car had no kinetic energy. Then it achieved kinetic energy equal to the work produced by the engine:


"KE=W=\frac{1}{2}mu^2."


At this point it's time to convert km/h to SI units m/s:


"u=v\cdot\frac{1000\text{ m/km}}{3600\text{ s/h}}=100\cdot\frac{1000\text{ m/km}}{3600\text{ s/h}}=27.78\text{ m/s}."

The average power output is how fast the work can be done:


"P=\frac{W}{t}=\frac{mu^2}{2t}=\frac{160\cdot27.78^2}{2\cdot5.5}=112953 \text{ W}."